Customers of a phone company can choose between two service plans for long distance calls. The first plan has an $11 monthly fee and charges an additional $0.14 for each minute of calls. The second plan has a $20 monthly fee and charges an additional $0.10 for each minute of calls. For how many minutes of calls will the costs of the two plans be equal
Answer by jim_thompson5910(28717)

First plan: y = 0.14x+11

Second plan: y = 0.10x+20


y = 0.14x+11

0.10x+20 = 0.14x+11


0.10x-0.14x = 11 - 20


-0.04x = -9


x = -9/(-0.04)


x = 225


So the two plans are equal when 225 minutes are used.